OpenStudy (anonymous):

f(x)/k(x) if f(x)=2x^3 and k(x)=x^-2/3

OpenStudy (anonymous):

k(x)=1/x^2 3

OpenStudy (anonymous):



OpenStudy (anonymous):

can you draw the answer please?

OpenStudy (campbell_st):

index law for division is subtract the powers of the same ponumeral \[2x^3/x^{-2/3} = 2x^{ 3 -(-2/3)} = 2x^{11/3}\]

OpenStudy (anonymous):

whats the domain?

OpenStudy (anonymous):

like (-infinity, infinity) and stuff

OpenStudy (anonymous):